8. (15 points) Let an = 5an-1 - 6an-2 with ao = 2 and a1 = 5. Let f:N → Z be given by f(n) = an. Prove that f(n) is (3).
Answer #1

first lets solve recurrence relation is an = 5an-1-600-2 so, 2-58+6 =0 D of g +62-18-ze No=(8-1) 2- (E-eje = (E-e (3-1) P so,

so, a an= a(z) + B(3) 2 = 2(2) + B (330 2 2 + B a, 5 = 2(2) + B (3) = 20 + 3ß d+B=2 > 204+ 2B = 4 /2 + 38=5 -ß =-) Ba| x

an = 2h+3h VI <3n+3n = 23 Since an is a c3 for some constant c then by definition of Big-O we can say that an is 0 (34)